OpenStudy (anonymous):

I need help finding the inverse of F(x)=5-8x/7x?

OpenStudy (tkhunny):

Please work on your notation. You have written \[F(x) = 5 - \frac{8x}{7x}\] I'm guessing you intended \[F(x) = \frac{5-8x}{7x}\] which would be written F(x) = (5-8x)/(7x). The usual swap and solve. Where are you getting stuck? \[x = \frac{5-8y}{7y}\] Solve for 'y'.